# Semiclassical Canovaccio for Composite Operators

### v1

@@ 1 Introduction
Conformal field theories (CFT)s describe a plethora of phase transitions, as established in pioneering works on criticality and the renormalization group Landau:1937obd ; Kadanoff:1966wm ; Kadanoff:1969zz ; Wilson:1973jj ; Wilson:1974mb . Significant progress has been made in two-dimensional and supersymmetric CFTs, and a variety of methodologies, such as the large quantum number expansion Hellerman:2015nra ; Komargodski:2012ek , conformal bootstrap Rattazzi:2008pe , as well as numerical simulations, have been developed to tackle different dynamical sectors of a wide class of CFTs. One major hurdle is constituted by the evaluation of scaling dimensions of generic neutral composite operators built out of several fields. In fact, these operators are notoriously hard to access even with the conformal bootstrap approach, which works best for the low-lying spectrum. Even in weakly-coupled theories, perturbation theory breaks down when determining correlators of these operators Goldberg:1990qk ; Badel:2019oxl . Additionally, operator mixing under renormalization group (RG) flow leads to anomalous dimension matrices whose size increases rapidly with the number of fields constituting the composite operators. This fact hampers the determination of the physical spectrum. Neutral composite operators are also relevant beyond CFTs as they appear in effective field theories such as the one employed to investigate new physics beyond the Standard Model (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri  [^1] We refer the reader interested in the multi-loop renormalization of the SMEFT to Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga for the complete $1$-loop renormalization and to Bern:2020ikv ; Aebischer:2022anv ; Jenkins:2023bls ; DiNoi:2024ajj ; Duhr:2025zqw ; Born:2024mgz for recent $2$-loop results.. General properties of the anomalous dimension matrix for a wide class of theories have been explored in e.g. Craigie:1983fb ; Alonso:2014rga ; Elias-Miro:2014eia ; Cheung:2015aba ; Craig:2019wmo ; Bern:2019wie ; Jiang:2020rwz ; Cao:2021cdt ; Cao:2023adc .
@@
where $C_{0}$ has been first computed in Antipin:2024ekk for the critical $\phi^{3}$, $\phi^{4}$, and $\phi^{6}$ theories. There, it was shown that in the small $\lambda n$ limit we recover the known perturbative results while predicting an infinite series of higher order terms. To this order in semiclassics the results depend purely on $n$, implying that the full spectrum of operators is degenerate. We further unveiled that the generic large $\lambda n$ behavior assumes the form $\Delta_{n,q_{\ell}}\propto n^{d/(d-1)}$. The $C_{1}(\lambda n,q_{\ell})$ term was presented first in Antipin:2025ilv for the Ising $\phi^{4}$ CFT. In the same work, within the semiclassical expansion, the dependence on $q_{\ell}$ first appeared showing how the spectrum degeneracy is lifted at NLO. One can view the results as a novel quantum field theory (QFT) application of the Gutzwiller trace formula Gutzwiller:1971fy , which expresses the energy spectrum as a trace over the periodic orbits and plays a key role in the semiclassical understanding of quantum chaos.
@@ 2 Free field theory
where $t_{0}$ is a zero-mode of the solution stemming from time-translation symmetry (the free parameter in Eq. (5)) and
@@ 3 The canovaccio as semiclassical blueprint
The introduction of the interaction substantially affects the spectrum of singlet composite operators via operator mixing. Let us consider spin-$s$ primary operators in the singlet representation of $\mathcal{G}$, which we can schematically denote as $\mathcal{O}_{s,p,n,i}=\partial^{s}\Box^{p}\phi^{n}$. With this notation, we intend the appropriate linear combination of terms with $n$ fields and $2p$ derivatives with fully contracted $\mathcal{G}$ and Lorentz indices and $s$ uncontracted derivatives such that $\mathcal{O}_{s,p,n,i}$ is a conformal primary. In other words, $\mathcal{O}_{s,p,n,i}$ is a primary eigenvector of the anomalous dimension matrix arising from the mixing of operators with the same spin $s$ and the same classical dimension $s+2p+n$. The additional index $i$ distinguishes between different primaries with the same schematic form, i.e., the same values of $s$, $p$, and $n$. In what follows, we will focus on primary operators transforming according to the traceless symmetric representation of the Lorentz group which can be labeled by $s$. As mentioned in the introduction, we denote their scaling dimensions as $\Delta_{n,q_{\ell}}$, where the precise meaning of this notation will be explained in Sec. 4. In the perturbative $\epsilon$-expansion $\Delta_{n,q_{\ell}}$ takes the form
@@
Our general strategy for the semiclassical evaluation of $\Delta_{n,q_{\ell}}$ consists of starting from homogeneous time-periodic classical solutions of the interacting theory on the cylinder. As long as the interaction strength is small, these periodic configurations can be seen as a perturbative deformation of the free field theory solution (9), and, therefore, we can expect them to be still related, via operator-state correspondence, to a primary operator of the CFT. We can then semiclassically compute the corresponding energy spectrum $E_{n,q_{\ell}}=r\Delta_{n,q_{\ell}}$ and validate the approach via explicit comparison with the $\epsilon$-expansion. On general grounds, these solutions assume an instantonic-like form $v(t)\sim{\epsilon^{-\beta}}\,\tilde{v}(t,m)$, for some positive $\beta$, where $m$ is a parameter to be fixed via the Bohr-Sommerfeld condition in Eq. (13)  [^2] While this form of the solution is correct for the simple examples considered in this work, multiple parameters may appear in more involved cases. Nevertheless, no changes in the scaling argument are expected.. The corresponding saddle point action scales as $\epsilon^{-2\beta}$, which is naturally the loop counting parameter of the semiclassical expansion. Moreover, since the action variable $I$ is quadratic in the fields, the Bohr-Sommerfeld condition can be seen as an equation determining $m=m(\epsilon^{2\beta}n)$. We are then led to consider the double scaling limit
@@
Accordingly, the semiclassical expansion is ultimately controlled by inverse powers of $n$ where the latter should be interpreted as the number of fields entering in the definition of the operator in the free field theory limit [^3] The underlying assumption is that, for every local operator for the free field theory obtained by setting $\kappa_{i}=0\ \forall\ i$ in $d$ dimensions, there exists a local operator of the interacting CFT in $d-\epsilon$ dimensions, which reduces to the former in the $\epsilon\to 0$ limit.. Consequently, the semiclassical expansion of $\Delta_{n,q_{\ell}}$ takes the form anticipated in Eq. (2)
@@
It follows that the procedure of tuning the theory to a Wilson-Fisher fixed point can also be seen as a trick to employ the power of conformal symmetry even away from conformality. As first elucidated in Antipin:2020rdw , for these purposes, the Wilson-Fisher fixed point can either be achieved for real or complex values of the couplings, and therefore it does not matter whether the resulting CFT is unitary or not. Note that for the anomalous dimension of operators in nontrivial representations of $\mathcal{G}$, this procedure has been applied to several QFTs, including the Standard Model Antipin:2023tar . It is interesting to observe that, while the anomalous dimensions are scheme-dependent quantities, the terms with the leading power of $n$ at any loop order are not, since they all stem from the classical calculation of $C_{0}$. Analogously, the terms with the subleading powers of $n$ coming from the one-loop semiclassical term $C_{1}$ are scheme independent as long as the one-loop renormalization constants are scheme independent as well.
@@
  • While we here focus on operators that are singlets of all the continuous symmetries of the theory, they can still transform in a nontrivial representation of a discrete symmetry group. For instance, in $Z_{2}$ invariant theories, a single periodic homogeneous saddle point, spontaneously breaking $Z_{2}$ describes both $Z_{2}$-even and $Z_{2}$-odd operators Antipin:2025ilv . However, the presence of a more involved discrete symmetry group may lead to the presence of multiple saddle points with different symmetry properties. Multiple saddle points generically emerge also when the scalar sector of the QFT contains scalar fields transforming according to different representations of $\mathcal{G}$.
@@
  • One may consider different scaling limits such as $N\to\infty$ with $n/N$ fixed in $O(N)$-invariant CFTs or $(N_{f}-N_{f}^{\rm crit})\to 0$ with fixed $(N_{f}-N_{f}^{\rm crit})n$ for fixed points of the Banks–Zaks type. For large charge operators, these directions have been explored in e.g. Alvarez-Gaume:2019biu ; Antipin:2022naw .
@@ 4 Semiclassical quantization of periodic solutions in QFT
The two eigenvalues of $\mathcal{M}$ can be written as $e^{\pm i\nu_{\ell}}$ where the $\nu_{\ell}$ variables are known as stability angles. A given periodic orbit is stable if and only if the stability angles are real. In what follows, we will assume a stable orbit and postpone the discussion of instabilities to Sec. 5.6. While it is clear that only $\nu_{\ell}\ \text{mod}~2\pi$ is physical, we here adopt conventions according to which the stability angles associated with the zero modes of $\mathcal{O}^{(2)}$ vanish. Remarkably, the determinant of $\mathcal{O}^{(2)}$ admits a simple expression in terms of the stability angles. In fact, applying the Gel’fand-Yaglom method, we obtain
@@
which determines ${\cal T}_{\rm saddle}={\cal T}_{\rm saddle}(E)$. Taking into account that one can traverse the basic periodic orbit $k$ times, one arrives at
@@
$C_{1}=\delta E_{1}+\frac{1}{{\cal T}}\sum_{\nu_{\ell}>0}\left(q_{\ell}+\frac{1}{2}\right)\nu_{\ell}\ .$ (47)
@@ 5.1 Classical solution and leading order: $C_{0}$
$nC_{0}=E_{\rm cl}=\Omega_{d-1}\braket{T_{00}}\rvert_{d=4}=\frac{2\pi^{2}m\,(1-m)}{\lambda\,(1-2m)^{2}}\ ,$ (55)
where, because the classical theory is conformal, we can fix $d=4$. The parameter $m$ is a function of $\lambda n$ obtained by inverting the Bohr-Sommerfeld condition (13), where
@@ 5.2 Renormalization of the action
In this section, we differentiate between bare and renormalized couplings and denote the former by $\lambda_{0}$. Our goal is to renormalize the exponent of Eq. (33), i.e. $\mathcal{S}_{\rm cl}-\frac{1}{2}\sum_{\nu_{\ell}>0}\nu_{\ell}$, employing dimensional regularization. As mentioned, when renormalizing the quartic coupling  [^5] Note that the fields do not get renormalized at the one-loop level. in the classical action, one isolates the $1/\epsilon$ pole needed to cancel the one-loop UV divergences arising when summing over the stability angles. The bare classical action in $d=4-\epsilon$ reads
$\mathcal{S}_{\rm cl}(\lambda_{0})=\Omega_{d-1}\int_{0}^{\cal T}\,dt\,\left(\frac{1}{2}\left(\frac{dv}{dt}\right)^{2}-\frac{\mu^{2}}{2}v^{2}-\frac{\lambda_{0}}{4}v^{4}\right)=-\frac{\pi^{2-\frac{\epsilon}{2}}(\epsilon-2)^{3}r^{-\epsilon}s(m)}{\lambda_{0}\Gamma\left(2-\frac{\epsilon}{2}\right)}\,$, (61)
@@
The order $\mathcal{O}\left({\lambda^{0}}\right)$ term is a quantum correction and in the $\epsilon\to 0$ limit, it reduces to
@@
$\frac{\partial\tilde{\mathcal{S}}(\lambda)}{\partial{\cal T}}\Big\rvert_{\lambda=\lambda_{*}}=\frac{\partial\tilde{\mathcal{S}}(\lambda)}{\partial m}\left(\frac{\partial{\cal T}}{\partial m}\right)^{-1}\Big\rvert_{\lambda=\lambda_{*}}=nC_{0}\left(\lambda_{*}n\right)-\frac{\lambda_{*}n}{2}\beta_{0}C_{0}\left(\lambda_{*}n\right)+\mathcal{O}\left({1/n}\right)\ ,$ (67)
@@
$\delta E_{1}=-\frac{\lambda_{*}n}{2}\beta_{0}C_{0}\left(\lambda_{*}n\right)\ .$ (68)

### v2

@@ 1 Introduction
Conformal field theories (CFT)s describe a plethora of phase transitions, as established in pioneering works on criticality and the renormalization group Landau:1937obd ; Kadanoff:1966wm ; Kadanoff:1969zz ; Wilson:1973jj ; Wilson:1974mb . Significant progress has been made in two-dimensional and supersymmetric CFTs, and a variety of methodologies, such as the large quantum number expansion Hellerman:2015nra ; Komargodski:2012ek , conformal bootstrap Rattazzi:2008pe , as well as numerical simulations, have been developed to tackle different dynamical sectors of a wide class of CFTs. One major hurdle is constituted by the evaluation of scaling dimensions of generic neutral composite operators built out of several fields. In fact, these operators are notoriously hard to access even with the conformal bootstrap approach, which works best for the low-lying spectrum. Even in weakly-coupled theories, perturbation theory breaks down when determining correlators of these operators Goldberg:1990qk ; Badel:2019oxl . Additionally, operator mixing under renormalization group (RG) flow leads to anomalous dimension matrices whose size increases rapidly with the number of fields constituting the composite operators. This fact hampers the determination of the physical spectrum. Neutral composite operators are also relevant beyond CFTs as they appear in effective field theories such as the one employed to investigate new physics beyond the Standard Model (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri  [^1] We refer the reader interested in the multi-loop renormalization of the SMEFT to Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga for the complete $1$-loop renormalization and to Bern:2020ikv ; Aebischer:2022anv ; Jenkins:2023bls ; DiNoi:2024ajj ; Duhr:2025zqw ; Born:2024mgz for recent $2$-loop results.. General properties of the anomalous dimension matrix for a wide class of theories have been explored in e.g., Craigie:1983fb ; Alonso:2014rga ; Elias-Miro:2014eia ; Cheung:2015aba ; Craig:2019wmo ; Bern:2019wie ; Jiang:2020rwz ; Cao:2021cdt ; Cao:2023adc .
@@
where $C_{0}$ has been first computed in Antipin:2024ekk for the critical $\phi^{3}$, $\phi^{4}$, and $\phi^{6}$ theories. There, it was shown that in the small $\lambda n$ limit we recover the known perturbative results while predicting an infinite series of higher order terms. To this order in semiclassics the results depend purely on $n$, implying that the full spectrum of operators is degenerate. We further unveiled that the generic large $\lambda n$ behavior assumes the form $\Delta_{n,q_{\ell}}\propto n^{d/(d-1)}$. The $C_{1}(\lambda n,q_{\ell})$ term was presented first in Antipin:2025ilv , for the Ising $\phi^{4}$ CFT. In the same work, within the semiclassical expansion, the dependence on $q_{\ell}$ first appeared showing how the spectrum degeneracy is lifted at NLO. One can view these results as a novel quantum field theory (QFT) application of the Gutzwiller trace formula Gutzwiller:1971fy , which expresses the energy spectrum as a trace over the periodic orbits and plays a key role in the semiclassical understanding of quantum chaos.
@@ 2 Free field theory
where $t_{0}$ is a zero-mode of the solution stemming from time-translation symmetry (the free parameters in Eq. (5)) and
@@ 3 The canovaccio as semiclassical blueprint
In the free field theory limit defined by $\epsilon\to 0$ and $\{\kappa_{i}^{*}(\epsilon)\}\to 0$, we can build a generic primary operator from components of the field and derivatives. Hence, a primary operator in the singlet representation of $\mathcal{G}$ can be schematically denoted as
$\mathcal{O}_{s,p,n,i}=\partial^{s}\Box^{p}\phi^{n}\ .$ (15)
With this notation, we intend the appropriate linear combination of terms with $n$ fields and $2p$ derivatives with fully contracted $\mathcal{G}$ and Lorentz indices and $s$ uncontracted derivatives such that $\mathcal{O}_{s,p,n,i}$ is a conformal primary. The additional index $i$ distinguishes between different primaries with the same schematic form, i.e., the same values of $s$, $p$, and $n$. The introduction of the interaction substantially affects the spectrum of singlet composite operators via operator mixing. That is, bare operators of the form of Eq. (15) mix under renormalization. Henceforth, a primary operator is no longer a single operator of the form of Eq. (15), but a linear combination of renormalized operators that diagonalize the anomalous dimension matrix. We here follow the standard convention of schematically denoting the interacting conformal primaries by those to which they reduce in the corresponding free field theory limit in $d$ dimensions  [^2] The underlying assumption is that, for every local operator for the free field theory obtained by setting $\kappa_{i}=0\ \forall\ i$ in $d$ dimensions, there exists a local operator of the interacting CFT in $d-\epsilon$ dimensions, which reduces to the former in the $\epsilon\to 0$ limit.. In what follows, we will focus on primary operators transforming according to the traceless symmetric representation of the Lorentz group, which can be labeled by $s$. As mentioned in the introduction, we denote their scaling dimensions as $\Delta_{n,q_{\ell}}$, where the precise meaning of this notation will be explained in Sec. 4. In the perturbative $\epsilon$-expansion $\Delta_{n,q_{\ell}}$ takes the form
@@
Our general strategy for the semiclassical evaluation of $\Delta_{n,q_{\ell}}$ consists of starting from homogeneous time-periodic classical solutions of the interacting theory on the cylinder. As long as the interaction strength is small, these periodic configurations can be seen as a perturbative deformation of the free field theory solution (9), and, therefore, we can expect them to be still related, via operator-state correspondence, to a primary operator of the CFT. We can then semiclassically compute the corresponding energy spectrum $E_{n,q_{\ell}}=r\Delta_{n,q_{\ell}}$ and validate the approach via explicit comparison with the $\epsilon$-expansion. On general grounds, these solutions assume an instantonic-like form $v(t)\sim{\epsilon^{-\beta}}\,\tilde{v}(t,m)$, for some positive $\beta$, where $m$ is a parameter to be fixed via the Bohr-Sommerfeld condition in Eq. (13)  [^3] While this form of the solution is correct for the simple examples considered in this work, multiple parameters may appear in more involved cases. Nevertheless, no changes in our scaling argument are expected.. The corresponding saddle point action scales as $\epsilon^{-2\beta}$, which is naturally the loop counting parameter of the semiclassical expansion. Moreover, since the action variable $I$ is quadratic in the fields, the Bohr-Sommerfeld condition can be seen as an equation determining $m=m(\epsilon^{2\beta}n)$. We are then led to consider the double scaling limit
@@
Accordingly, the semiclassical expansion is ultimately controlled by inverse powers of $n$ where, as mentioned, the latter should be interpreted as the number of fields entering in the definition of the operator in the free field theory limit. Consequently, the semiclassical expansion of $\Delta_{n,q_{\ell}}$ takes the form anticipated in Eq. (2)
@@
It follows that the procedure of tuning the theory to a Wilson-Fisher fixed point can also be seen as a trick to employ the power of conformal symmetry even away from conformality. As first elucidated in Antipin:2020rdw , for these purposes, the Wilson-Fisher fixed point can either be achieved for real or complex values of the couplings, and therefore it does not matter whether the resulting CFT is unitary or not. Note that for the anomalous dimension of operators in nontrivial representations of $\mathcal{G}$, this procedure has been applied to several QFTs including the Standard Model Antipin:2023tar . It is interesting to observe that, while the anomalous dimensions are scheme-dependent quantities, the terms with the leading power of $n$ at any loop order are not, since they all stem from the classical calculation of $C_{0}$. Analogously, the terms with the subleading powers of $n$ coming from the one-loop semiclassical term $C_{1}$ are scheme independent as long as the one-loop renormalization constants are scheme independent as well.
@@
  • While we here focus on operators that are singlets of all the continuous symmetries of the theory, they can still transform in a nontrivial representation of a discrete symmetry group. For instance, in $Z_{2}$ invariant theories, a single periodic homogeneous saddle point spontaneously breaking $Z_{2}$ describes both $Z_{2}$-even and $Z_{2}$-odd operators Antipin:2025ilv . However, the presence of a more involved discrete symmetry group may lead to the presence of multiple saddle points with different symmetry properties. Multiple saddle points generically emerge also when the scalar sector of the QFT contains scalar fields transforming according to different representations of $\mathcal{G}$.
@@
  • One may consider different scaling limits such as $N\to\infty$ with $n/N$ fixed in $O(N)$-invariant CFTs or $(N_{f}-N_{f}^{\rm crit})\to 0$ with fixed $(N_{f}-N_{f}^{\rm crit})n$ for fixed points of the Banks–Zaks type. For large charge operators, these directions have been explored in e.g., Alvarez-Gaume:2019biu ; Antipin:2022naw .
@@ 4 Semiclassical quantization of periodic solutions in QFT
The two eigenvalues of $\mathcal{M}$ can be written as $e^{\pm i\nu_{\ell}}$ where the $\nu_{\ell}$ variables are known as stability angles. A given periodic orbit is stable if and only if the stability angles are real. In what follows, we will assume a stable orbit and postpone the discussion of instabilities to Sec. 5.6. While it is clear that only $\nu_{\ell}\ \text{mod}~2\pi$ is physical, we here adopt conventions according to which the stability angles associated with the zero modes of $\mathcal{O}^{(2)}$ vanish. Remarkably, the determinant of $\mathcal{O}^{(2)}$ admits a simple expression in terms of the stability angles. In fact, applying the Gel’fand-Yaglom method, we obtain Dunne:2007rt
@@
which determines ${\cal T}_{\rm saddle}={\cal T}_{\rm saddle}(E)$, thereby fixing $\cal T$} as the period of the classical orbit with the given energy $E$. Taking into account that one can traverse the basic periodic orbit $k$ times, one arrives at
@@
$nC_{0}=rE_{cl}\,,\qquad C_{1}=r\delta E_{1}+\frac{r}{{\cal T}}\sum_{\nu_{\ell}>0}\left(q_{\ell}+\frac{1}{2}\right)\nu_{\ell}\ .$ (48)
@@
As evident from the formalism, the generalization to multiple fields involves the presence of a functional determinant for each field, which in turn leads to multiple sets of stability angles and associated occupation numbers $q_{\ell}$. This will indeed be the case for the $O(N)$ scalar theories investigated in the next sections.
@@ 5.1 Classical solution and leading order: $C_{0}$
$nC_{0}=rE_{\rm cl}=\Omega_{d-1}r^{d}\braket{T_{00}}\rvert_{d=4}=\frac{2\pi^{2}m\,(1-m)}{\lambda\,(1-2m)^{2}}\ ,$ (56)
where $T_{\mu\nu}$ is the stress energy tensor and, because the classical theory is conformal, we can fix $d=4$. The parameter $m$ is a function of $\lambda n$ obtained by inverting the Bohr-Sommerfeld condition (13), where
@@ 5.2 Renormalization of the action
In this section, we differentiate between bare and renormalized couplings and denote the former by $\lambda_{0}$. Our goal is to renormalize the exponent of Eq. (34), i.e., $\mathcal{S}_{\rm cl}-\frac{1}{2}\sum_{\nu_{\ell}>0}\nu_{\ell}$, employing dimensional regularization. As mentioned, when renormalizing the quartic coupling  [^5] Note that the fields do not get renormalized at the one-loop level. in the classical action, one isolates the $1/\epsilon$ pole needed to cancel the one-loop UV divergences arising when summing over the stability angles. The bare classical action in $d=4-\epsilon$ reads
$\mathcal{S}_{\rm cl}(\lambda_{0})=\Omega_{d-1}r^{d-1}\int_{0}^{\cal T}\,dt\,\left(\frac{1}{2}\left(\frac{dv}{dt}\right)^{2}-\frac{\mu^{2}}{2}v^{2}-\frac{\lambda_{0}}{4}v^{4}\right)=-\frac{\pi^{2-\frac{\epsilon}{2}}(\epsilon-2)^{3}r^{-\epsilon}s(m)}{\lambda_{0}\Gamma\left(2-\frac{\epsilon}{2}\right)}\,$, (62)
@@
The order $\mathcal{O}\left({\lambda^{0}}\right)$ term is a quantum correction and in the $\epsilon\to 0$ limit it reduces to
@@
$r\frac{\partial\tilde{\mathcal{S}}(\lambda)}{\partial{\cal T}}\Big\rvert_{\lambda=\lambda_{*}}=r\frac{\partial\tilde{\mathcal{S}}(\lambda)}{\partial m}\left(\frac{\partial{\cal T}}{\partial m}\right)^{-1}\Big\rvert_{\lambda=\lambda_{*}}=nC_{0}\left(\lambda_{*}n\right)-\frac{\lambda_{*}n}{2}\beta_{0}C_{0}\left(\lambda_{*}n\right)+\mathcal{O}\left({1/n}\right)\ ,$ (68)
@@
$r\delta E_{1}=-\frac{\lambda_{*}n}{2}\beta_{0}C_{0}\left(\lambda_{*}n\right)\ .$ (69)